OpenStudy (anonymous):

I WILL GIVE MEDAL AND BECOME A FAN 5x + 3y = -2 Isolate y: ______ on both sides ______ on both sides 5x + 3y =-2 3y = ______ y = ______ The rate of change is _____. Please explain this to me simply, thank you very much! :)

OpenStudy (anonymous):

im sorry but i dont think i would be much help here i know last semester i had a problem similar and doing it the same way by putting it in slope intercept or F(x) form you would get: 3y = 5x -(-2) 3y = 5x + 2 y = 5x + 2 / 3

OpenStudy (anonymous):

Thanks for the help, do you know how I can fill this? Isolate y: ______ on both sides ______ on both sides 5x + 3y =-2 3y = ______ y = ______ The rate of change is _____.

OpenStudy (anonymous):

\[\frac{ 5x + 2 }{ 3 }\] it should look like this if i am correct